You have three solutions, “A”, “B” and “C”. Each must be diluted by the following dilution factors: 4X, 10X and 50X respectively. What would be the final volume of a solution prepared by adding each of the ingredients to 100 mL of solvent?
Answer #1

Now for each three solutions namely A,B and C we start with 100 ml.of solution.

For A we have to do 4X that is 4 times which is 4 × 100 = 400 ml

For B similarly it is 10 X = 10 × 100 = 1000ml

For C similarly it is 50X = 50 × 100= 5000ml